# A Two-Stage Pretraining-Finetuning Framework for Treatment Effect Estimation with Unmeasured Confounding

### v1

@@ 4.1.2. Reconstruction and Prediction
We then use the representations $Z$ to estimate the potential outcomes with two $l_{p}-$layer prediction heads $h_{0}$ and $h_{1}$, which are the predictors for control and treatment outcomes, respectively. Note that unmeasured confounding in the observational data can lead to biased estimations of potential outcomes, we refer to the prediction result $\tilde{Y}^{0}=h_{0}(Z)=h_{0}(\phi(X))$ as the pseudo control outcome and $\tilde{Y}^{1}=h_{1}(Z)=h_{1}(\phi(X))$ as the pseudo treatment outcome. To enhance comparability between the treatment group and control group, we employ a reweighting technique to balance the two groups. Formally, let $f_{h}(x,t)=h_{t}(\phi(x))$ with $t\in\{0,1\}$ be the predicted potential outcomes by via the two heads $h_{0}$ and $h_{1}$, the loss for outcome prediction is as follows:
@@
$\min_{\phi,\psi,h_{0},h_{1}}\mathcal{L}_{f}+\lambda_{1}\mathcal{L}_{rec}+\lambda_{2}\mathcal{L}_{imb},$
@@ 4.2.2. Prediction
Similarly to the first stage, we design two $l_{p}-$layer prediction heads $g_{0}$ and $g_{1}$ to estimate the potential outcomes under control and treatment groups with unmeasured confoundings, respectively. Notice that $g_{0},g_{1}$ and $h_{0},h_{1}$ have the same depth $l_{p}$, yet every layer of $g_{t}$ has a larger or equal width than $h_{t}$. For $t\in\{0,1\}$, we define the header $h_{t}$ as:
@@
with $W_{g_{t}}^{(l)},b_{g_{t}}^{(l)},a_{g_{t}}^{(l)}$ having similar meanings as $W_{h_{t}}^{(l)},b_{h_{t}}^{(l)},a_{h_{t}}^{(l)}$ but for $g_{t}$, $\hat{Y}_{t}$ is the final prediction for potential outcome under treatment $t$. In our design, every layer of $g_{t}$ has a larger or equal width than $h_{t}$, thus the dimension of $W_{g_{t}}^{(l)},b_{g_{t}}^{(l)},a_{g_{t}}^{(l)}$ is no less than that of $W_{h_{t}}^{(l)},b_{h_{t}}^{(l)},a_{h_{t}}^{(l)}$ respectively. We divide the parameters of $g_{t}$ as:
@@
Initialization.
As in the first stage, we denote $f_{g}(x,t)=g_{t}([\phi(x)^{\intercal}|\phi^{U}(x)^{\intercal}]^{\intercal})$ for $t\in\{0,1\}$, where $[\phi(x)^{\intercal}|\phi^{U}(x)^{\intercal}]$ is the concatenated covariate representation. Given the RCT data $\{(X_{i},T_{i},Y_{i},G_{i}=1):i\in\mathcal{A}\}$, the prediction loss is computed as:
Regularization.
The presence of unmeasured confounding may cause a slight shift in the distribution of OBS data from the RCT data. Therefore the second-stage fine-tuned model should not deviate significantly from the first-stage model. We denote the initial value of $\theta_{g_{t}}$ as $\theta_{g_{t}}^{0}$. In order to constrain the deviation from the initial value, we include an $l_{2}-$norm in the loss function:
Overall, the training objective of the second stage is given by:
where $\lambda_{3}>0$ and $\lambda_{4}>0$ are tunable hyperparameters. Note that during the second-phase training, we froze the parameters of the representation module $\phi$ and the decoder network $\psi$, while train the representation adapter module $\phi^{U}$ and the two prediction heads $g_{0},g_{1}$. We summarize the whole learning algorithm in Alg. 1.
Compared to residual correction methods as in Kallus et al. (Kallus et al., 2018), our representation adapter module guarantees a stronger representation ability, relaxing the linearly additive assumption. While compared to methods that jointly learn residual and CATE, one advantage is that when RCT data are limited, our proposed partial initialization strategy in the TSPF framework can avoid overfitting.
@@ 5.1. Datasets
Following previous studies (Shalit et al., 2017; Louizos et al., 2017; Yoon et al., 2018), we conduct experiments on two publicly available datasets, namely IHDP (Hill, 2011) and Jobs (Shalit et al., 2017). The IHDP is a semi-synthetic dataset for causal effect estimation. The dataset is based on the Infant Health and Development Program, where the covariates are obtained by a randomized experiment investigating the effect of home visits by specialists on future cognitive scores. It consists of 747 units (19% treated, 81% control) and 25 covariates measuring the children and their mothers. The Jobs is a common benchmark dataset developed by LaLonde in 1986, studying the change of income and employment status after job training. We use an extended version of Jobs that comprises about 3,000 units (10% treated, 90% control) with 17 covariates.
@@ 5.2. Data Preprocessing
For both IHDP and Jobs, we simulate unmeasured confounding by generating a $c$-dimensional confounder $U_{i}\in\mathbb{R}^{c}$. To make sure the $U_{i}$ has a non-zero effect on $Y_{i}$ and $T_{i}$, we generate the data below:
@@
where $\mathcal{N}(\mu,D)$ denotes the normal distribution with mean $\mu$ and variance $D$, $\mathcal{U}(a,b)$ is the uniform distribution on interval $(a,b)$, $\text{Bern}(p)$ means the Bernoulli distribution with probability $p$, $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function. Note that we keep $E[Y_{i}^{1}]=E[Y_{i}^{0}]+4$ as the same as the IHDP dataset and we let $E[W_{i}]>0\text{ for }i\in\{2,4,6\}$ to ensure the non-zero effect of $U_{i}$. The unmeasured confounding strength parameter $c$ is set to 30. Then we slice the training, validation, and test sets in the ratio of 63/27/10. In addition, to obtain a separate RCT training dataset for data fusion, we first randomly split $r\%$ of the training samples, and then assign treatments $T^{new}_{i}$ according to the following formula and replace the factual treatment $T_{i}$ and outcome $Y^{f}_{i}$ to obtain a RCT dataset:
$T^{new}_{i}=\operatorname{Bern}(0.5),~{}~{}~{}Y^{new}_{i}=\mathbb{I}(T^{new}_{i}=T_{i})(Y^{f}_{i}-Y^{cf}_{i})+Y^{cf}_{i},$
where $\mathbb{I}(\cdot)$ is the indicator function, $Y^{f}_{i}=T_{i}Y_{i}^{1}+(1-T_{i})Y_{i}^{0}$ is the factual outcome, and $Y^{cf}_{i}=T_{i}Y_{i}^{0}+(1-T_{i})Y_{i}^{1}$ is the counterfactual outcome. The RCT data ratio $r\%$ is set to 10% unless otherwise stated. Finally, we replace the $T_{i}$ and $Y_{i}$ using the above formula for all samples in the validation set.
Table 2. Ablation study on the pretraining stage and fine-tuning stage, the initialization strategy, and the training strategy of freezing foundational representation module and the initialized parameters.
IHDP Jobs
In-sample Out-sample In-sample Out-sample
$\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$ $\sqrt{\epsilon_{\text{PEHE}}}$ $\epsilon_{\text{ATE}}$
TSFP w/o $\mathcal{L}_{rec}$ and $\mathcal{L}_{imb}$ 0.25 $\pm$ 0.05 0.07 $\pm$ 0.02 0.29 $\pm$ 0.08 0.07 $\pm$ 0.02 0.19 $\pm$ 0.03 0.04 $\pm$ 0.03 0.18 $\pm$ 0.02 0.04 $\pm$ 0.03
TSFP w/o $\mathcal{L}_{rec}$ $0.21\pm 0.04$ $0.05\pm 0.02$ $0.24\pm 0.05$ $0.05\pm 0.02$ $0.14\pm 0.01$ $0.03\pm 0.01$ $0.13\pm 0.02$ $0.03\pm 0.01$
TSFP w/o $\mathcal{L}_{imb}$ $0.18\pm 0.03$ $0.03\pm 0.02$ $0.19\pm 0.05$ $0.05\pm 0.02$ $0.12\pm 0.02$ $0.03\pm 0.02$ $0.10\pm 0.04$ $0.03\pm 0.01$
TSFP w/o $\mathcal{L}_{MI}$ and $\mathcal{L}_{shift}$ 0.23 $\pm$ 0.03 0.06 $\pm$ 0.01 0.26 $\pm$ 0.05 0.07 $\pm$ 0.02 0.16 $\pm$ 0.04 0.03 $\pm$ 0.01 0.18 $\pm$ 0.04 0.03 $\pm$ 0.01
TSFP w/o $\mathcal{L}_{MI}$ 0.17 $\pm$ 0.04 0.04 $\pm$ 0.01 0.18 $\pm$ 0.04 0.05 $\pm$ 0.02 0.12 $\pm$ 0.03 0.02 $\pm$ 0.01 0.15 $\pm$ 0.03 0.02 $\pm$ 0.01
TSFP w/o $\mathcal{L}_{shift}$ 0.19 $\pm$ 0.04 0.05 $\pm$ 0.01 0.21 $\pm$ 0.05 0.05 $\pm$ 0.02 0.15 $\pm$ 0.05 0.03 $\pm$ 0.01 0.16 $\pm$ 0.04 0.02 $\pm$ 0.01
TSFP not freezing representation $0.17\pm 0.02$ $0.03\pm 0.03$ $0.20\pm 0.05$ $0.05\pm 0.04$ 0.10 $\pm$ 0.03 0.01 $\pm$ 0.01 0.09 $\pm$ 0.03 0.01 $\pm$ 0.01
TSFP freezing initialized parameters $0.16\pm 0.08$ $0.03\pm 0.02$ $0.19\pm 0.05$ $0.04\pm 0.02$ 0.09 $\pm$ 0.03 0.01 $\pm$ 0.01 0.08 $\pm$ 0.04 0.01 $\pm$ 0.01
TSFP w/o initialization $0.35\pm 0.09$ $0.08\pm 0.07$ $0.38\pm 0.11$ $0.07\pm 0.06$ 0.42 $\pm$ 0.12 0.06 $\pm$ 0.04 0.41 $\pm$ 0.13 0.06 $\pm$ 0.04
TSFP w/o initialization and $\mathcal{L}_{shift}$ $0.43\pm 0.09$ $0.11\pm 0.07$ $0.45\pm 0.10$ $0.12\pm 0.06$ 0.54 $\pm$ 0.13 0.09 $\pm$ 0.05 0.53 $\pm$ 0.14 0.11 $\pm$ 0.05
TSFP w/o $Z^{U}$ $0.18\pm 0.03$ $0.05\pm 0.02$ $0.19\pm 0.04$ $0.05\pm 0.02$ 0.14 $\pm$ 0.03 0.02 $\pm$ 0.01 0.16 $\pm$ 0.04 0.02 $\pm$ 0.01
TSFP $0.13\pm 0.02$ $0.02\pm 0.02$ $0.16\pm 0.04$ $0.04\pm 0.02$ 0.09 $\pm$ 0.03 0.01 $\pm$ 0.01 0.06 $\pm$ 0.01 0.01 $\pm$ 0.01
Figure 2. The experiment results for the IHDP dataset under different RCT data ratios in the training data.
@@ 5.3.1. Baselines
  • DRCFR (Hassanpour and Greiner, 2019): DRCFR aims to learn disentangled representations and address selection bias in CATE estimation.
  • Twostep linear (Kallus et al., 2018): Twostep linear method uses OBS data to learn a biased estimate for the treatment effect and then aims to remove the bias using RCT data.
  • CorNet (Hatt et al., 2022b): CorNet uses the RCT data to learn a non-linear bias function in the second stage.
@@ 5.3.3. Experimental Details
We implement TSPF with a multi-layer perceptron (Haykin, 1994) with 2 layers for our representation and reconstruction modules as well as the prediction heads in both stages using the PyTorch framework. We tune the hyper-parameters in the loss functions from $51e-$ 5 to $0.1$, the learning rate from $51e-$ 5 to $21e-$ 2, and the weight decay parameter from $51e-$ 5 to $21e-$ 2. The batch size is set to 256 in all scenarios and the optimizer is selected in Adam and SGD. We report both in-sample and out-of-sample results for metrics $\sqrt{\epsilon_{\mathrm{PEHE}}}$ and $\epsilon_{\mathrm{ATE}}$ in our experiments.
@@ 5.4. Performance Analysis
Table 1 shows the prediction performance with varying baselines and our methods. First, representation-based methods generally outperform generation-based methods and meta-learners, which shows the effectiveness of causal representation learning. Note that our proposed TSPF exhibits the most competitive performance in most cases, outperforming all the baselines except for the $\epsilon_{\mathrm{ATE}}$ metric in out-sample scenario. In the comparison of two-step methods, we can see that CorNet (Hatt et al., 2022b), which adopts a nonlinear residual module, outperforms the Twostep linear method based on the assumption that the residual $\eta(X_{i})$ is linear and additive. In addition, our method stably outperforms the other two-stage methods, showing the effectiveness of our pretraining-finetuning framework.
5.5. Ablation Study
After evaluating the overall performance of our method, we perform ablation studies to validate the effectiveness of each module in our approach. Firstly, we consider employing all the modules in the first stage and taking away the mutual information term and/or the shift regularization term. Then we include all the modules in the second stage and test the effect of removing the reconstruction module and/or the IPM distance. To validate the soundness of our training algorithm, especially the initialization strategy, we also evaluate the performance of TSPF when not freezing the foundational representation module $\phi$, freezing the initialized parameters in $g_{0}$ and $g_{1}$, without the initialization strategy, and without the augmented representation $Z^{U}$. The results are presented in Table 2, which shows that each component in our framework plays an important role in accurately estimating CATE. In the first/second stage, $\mathcal{L}_{rec}$/$\mathcal{L}_{shift}$ is more important than $\mathcal{L}_{imb}$/$\mathcal{L}_{MI}$. In addition, our initialization strategy significantly improves the performance.
5.6. In-Depth Analysis
Table 3. Experimental results on the IHDP dataset with unmeasured confounding strength $c=10$ and $c=50$. The best result is bolded and the second best is underlined.
unmeasured confounding strength $c=10$ unmeasured confounding strength ${c=50}$
In-sample Out-sample In-sample Out-sample
T-learner 0.42 $\pm$ 0.06 0.03 $\pm$ 0.02 0.47 $\pm$ 0.08 0.06 $\pm$ 0.05 0.48 $\pm$ 0.05 0.05 $\pm$ 0.02 0.53 $\pm$ 0.06 0.04 $\pm$ 0.03
S-learner 0.99 $\pm$ 0.26 0.05 $\pm$ 0.04 1.29 $\pm$ 0.34 0.13 $\pm$ 0.11 1.17 $\pm$ 0.11 0.05 $\pm$ 0.03 1.51 $\pm$ 0.19 0.11 $\pm$ 0.10
DR-learner 0.69 $\pm$ 0.12 0.06 $\pm$ 0.03 0.72 $\pm$ 0.13 0.10 $\pm$ 0.05 0.73 $\pm$ 0.11 0.07 $\pm$ 0.04 0.80 $\pm$ 0.13 0.08 $\pm$ 0.06
Causal Forest 1.93 $\pm$ 0.31 0.08 $\pm$ 0.07 1.85 $\pm$ 0.28 0.24 $\pm$ 0.22 1.98 $\pm$ 0.18 0.13 $\pm$ 0.06 2.04 $\pm$ 0.34 0.29 $\pm$ 0.19
TARNet 0.26 $\pm$ 0.06 0.03 $\pm$ 0.02 0.27 $\pm$ 0.13 0.04 $\pm$ 0.04 0.52 $\pm$ 0.20 0.09 $\pm$ 0.09 0.53 $\pm$ 0.24 0.11 $\pm$ 0.11
CEVAE 3.09 $\pm$ 0.92 0.19 $\pm$ 0.16 3.19 $\pm$ 1.22 0.41 $\pm$ 0.33 3.50 $\pm$ 0.51 0.25 $\pm$ 0.16 3.39 $\pm$ 0.45 0.65 $\pm$ 0.32
SCIGAN 2.74 $\pm$ 0.68 0.62 $\pm$ 0.28 2.77 $\pm$ 0.89 0.56 $\pm$ 0.30 2.94 $\pm$ 0.42 0.46 $\pm$ 0.25 2.90 $\pm$ 0.35 0.75 $\pm$ 0.38
DragonNet 0.19 $\pm$ 0.04 0.03 $\pm$ 0.02 0.23 $\pm$ 0.05 0.06 $\pm$ 0.07 0.23 $\pm$ 0.05 0.06 $\pm$ 0.07 0.29 $\pm$ 0.13 0.07 $\pm$ 0.04
DESCN 0.36 $\pm$ 0.17 0.04 $\pm$ 0.03 0.42 $\pm$ 0.26 0.08 $\pm$ 0.09 0.46 $\pm$ 0.14 0.09 $\pm$ 0.05 0.55 $\pm$ 0.23 0.08 $\pm$ 0.06
DRCFR 0.70 $\pm$ 0.32 0.09 $\pm$ 0.08 0.92 $\pm$ 0.74 0.18 $\pm$ 0.22 1.21 $\pm$ 0.80 0.30 $\pm$ 0.20 1.17 $\pm$ 0.72 0.26 $\pm$ 0.18
Twostep linear 0.70 $\pm$ 0.16 0.29 $\pm$ 0.25 0.83 $\pm$ 0.30 0.30 $\pm$ 0.39 0.76 $\pm$ 0.21 0.42 $\pm$ 0.23 0.83 $\pm$ 0.25 0.33 $\pm$ 0.31
CorNet 0.23 $\pm$ 0.12 0.05 $\pm$ 0.04 0.27 $\pm$ 0.13 0.04 $\pm$ 0.04 0.39 $\pm$ 0.19 0.08 $\pm$ 0.08 0.38 $\pm$ 0.18 0.07 $\pm$ 0.07
TSFP (ours) 0.15 $\pm$ 0.04 0.02 $\pm$ 0.02 0.15 $\pm$ 0.04 0.02 $\pm$ 0.02 0.15 $\pm$ 0.03 0.02 $\pm$ 0.01 0.17 $\pm$ 0.07 0.02 $\pm$ 0.02
Figure 3. Sensitivity analysis on the reconstruction loss strength $\lambda_{1}$, covariate imbalancing loss strength $\lambda_{2}$, mutual information loss strength $\lambda_{3}$ and RCT shift loss strength $\lambda_{4}$ on the IHDP dataset.
5.6.1. RCT Data Ratio
To validate whether our framework performs well with limited RCT data, we further explore the effect of using different ratios of RCT data in the training set. We conduct experiments with a sequence of increasing ratios from 2% to 20%. For each scenario, the optimal network size and hyper-parameters are finetuned to ensure fair comparison. The results in Figure 2 show that our method is robust to the change of RCT data ratio and outperforms the baselines under different RCT ratios.
5.6.2. Unmeasured Confounding Strength
Our approach aims to achieve unbiased estimation of CATE in the presence of unmeasured confounding. To verify the validity under different strengths of unmeasured confounding, we add a comparison of the performance under low ($c=10$) and high ($c=50$) unmeasured confounding effect, which is shown in Table 3. The results show that our method stably outperforms baselines with varying $c$.
5.7. Sensitivity Analysis
Despite there are many hyper-parameters in our TSPF framework, they present in two separate stages, i.e, $\lambda_{1}$ and $\lambda_{2}$ in the first stage, while $\lambda_{3}$ and $\lambda_{4}$ in the second stage. As a result, the hyper-parameter search space is significantly reduced, compared with simultaneously searching four hyper-parameters. We explore the sensitivity of the hyper-parameters on the IHDP dataset and the results are shown in Figure 3. For all four hyper-parameters, the best result is achieved with moderate values. The performance dramatically drops with low values of hyper-parameters, which further shows the effectiveness of each component in our method.

### v2

@@ 3.1.2. Reconstruction and Prediction
We then use the representations $Z$ to estimate the potential outcomes with two $l_{p}-$layer prediction heads $h_{0}$ and $h_{1}$, which are the predictors for control and treatment outcomes, respectively. Note that hidden confounding in the observational data can lead to biased estimations of potential outcomes, we refer to the prediction result $\tilde{Y}^{0}=h_{0}(Z)=h_{0}(\phi(X))$ as the pseudo control outcome and $\tilde{Y}^{1}=h_{1}(Z)=h_{1}(\phi(X))$ as the pseudo treatment outcome. To enhance comparability between the treatment group and control group, we employ a reweighting technique to balance the two groups. Formally, let $f_{h}(x,t)=h_{t}(\phi(x))$ with $t\in\{0,1\}$ be the predicted potential outcomes by via the two heads $h_{0}$ and $h_{1}$, the loss for outcome prediction is as follows:
@@
$\min_{\phi,\psi,h_{0},h_{1}}\mathcal{L}_{f}+\lambda_{1}\mathcal{L}_{rec}+\lambda_{2}\mathcal{L}_{unb},$
@@ 3.2.2. Prediction
Similarly to the first stage, we design two $l_{p}-$layer prediction heads $g_{0}$ and $g_{1}$ to estimate the potential outcomes under control and treatment groups with hidden confoundings, respectively. Notice that $g_{0},g_{1}$ and $h_{0},h_{1}$ have the same depth $l_{p}$, yet every layer of $g_{t}$ has a larger or equal width than $h_{t}$. For $t\in\{0,1\}$, we define the header $h_{t}$ as:
@@
with $W_{g_{t}}^{(l)},b_{g_{t}}^{(l)},a_{g_{t}}^{(l)}$ having similar meanings as $W_{h_{t}}^{(l)},b_{h_{t}}^{(l)},a_{h_{t}}^{(l)}$ but for $g_{t}$, $\hat{Y}_{t}$ is the final prediction for potential outcome under treatment $t$. In our design, every layer of $g_{t}$ has a larger or equal width than $h_{t}$, thus the dimension of $W_{g_{t}}^{(l)},b_{g_{t}}^{(l)},a_{g_{t}}^{(l)}$ is no less than that of $W_{h_{t}}^{(l)},b_{h_{t}}^{(l)},a_{h_{t}}^{(l)}$, respectively. We divide the parameters of $g_{t}$ as:
@@
Initialization.
As in the first stage, we denote $f_{g}(x,t)=g_{t}([\phi(x)^{\intercal}|\phi^{U}(x)^{\intercal}]^{\intercal})$ for $t\in\{0,1\}$, where $[\phi(x)^{\intercal}|\phi^{U}(x)^{\intercal}]$ is the concatenated covariate representation. Given the RCT data $\{(X_{i},T_{i},Y_{i},G_{i}=1):i\in\mathcal{R}\}$, the prediction loss is computed as:
Regularization.
Since small-scale RCT data may cause overfitting during the finetuning phase, we further introduce a constraint to ensure that the parameters of the second-stage finetuned model do not deviate significantly from the parameters of the first-stage trained model. Denote the initial value of $\theta_{g_{t}}$ as $\theta_{g_{t}}^{0}$, to constrain the deviation from the initial value, we propose to use $l_{2}$-norms in the training loss objective:
Overall, the training loss of the second stage is given by:
where $\lambda_{3}$ and $\lambda_{4}$ are tunable hyperparameters. Note that during the second-phase training, we froze the parameters of the representation module $\phi$ and the decoder network $\psi$, while train the representation adapter module $\phi^{U}$ and the two prediction heads $g_{0},g_{1}$. We summarize the whole learning algorithm in Alg. 1.
@@
Compared to residual correction methods as in Kallus et al. (Kallus et al., 2018), our representation adapter module guarantees a stronger representation ability, relaxing the linearly additive assumption. More importantly, when RCT data are limited, our proposed partial initialization strategy in the TSPF framework can avoid overfitting.
@@ 4.1. Datasets
Following previous studies (Shalit et al., 2017; Louizos et al., 2017; Yoon et al., 2018), we conduct experiments on two semi-synthetic datasets, namely IHDP (Hill, 2011) and Jobs (Shalit et al., 2017). The IHDP is a semi-synthetic dataset for causal effect estimation. The dataset is based on the Infant Health and Development Program, where the covariates are obtained by a randomized experiment investigating the effect of home visits by specialists on future cognitive scores. It consists of 747 units (19% treated, 81% control) and 25 covariates measuring the children and their mothers. The Jobs is a common benchmark dataset developed by LaLonde in 1986, studying the change of income and employment status after job training. We use an extended version of Jobs that comprises about 3,000 units (10% treated, 90% control) with 17 covariates.
@@ 4.2. Data Preprocessing
For both IHDP and Jobs, we simulate hidden confounding by generating a $c$-dimensional confounder $U_{i}\in\mathbb{R}^{c}$. To make sure the $U_{i}$ has a non-zero effect on $Y_{i}$ and $T_{i}$, we generate the data below:
@@
where $\mathcal{N}(\mu,D)$ denotes the normal distribution with mean $\mu$ and variance $D$, $\mathcal{U}(a,b)$ is the uniform distribution on interval $(a,b)$, $\text{Bern}(p)$ means the Bernoulli distribution with probability $p$, $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function. Note that we keep $E[Y_{i}^{1}]=E[Y_{i}^{0}]+4$ as the same as the IHDP dataset and we let $E[W_{i}]>0,i\in\{2,4,6\}$ to ensure the non-zero effect of $U_{i}$. The hidden confounding strength parameter $c$ is set to 30 in our experiments. Then we slice the training, validation, and test sets in the ratio of 63/27/10. In addition, to obtain a separate RCT training dataset for data fusion, we first randomly split 10% of the training samples, and then assign treatments $T^{new}_{i}$ according to the following formula and replace the factual treatment $T_{i}$ and outcome $Y^{f}_{i}$ to obtain a RCT dataset:
$T^{new}_{i}=\operatorname{Bern}(0.5),~{}~{}~{}Y^{new}_{i}=\mathbb{I}\{T^{new}_{i}=T_{i}\}(Y^{f}_{i}-Y^{cf}_{i})+Y^{cf}_{i},$
where $\mathbb{I}$ is the indicator function, $Y^{f}_{i}=T_{i}Y_{i}^{1}+(1-T_{i})Y_{i}^{0}$ is the factual outcome, and $Y^{cf}_{i}=T_{i}Y_{i}^{0}+(1-T_{i})Y_{i}^{1}$ is the counterfactual outcome. Finally, we replace the treatment $T_{i}$ and factual outcome $Y_{i}$ using the above formula for all samples in the validation set.
@@ 4.3.1. Baselines
  • DRCFR (Hassanpour and Greiner, 2019): DRCFR aims to learn disentangled representations of covariates and address selection bias in CATE estimation.
  • CorNet (Hatt et al., 2022a): CorNet leverages OBS data to learn a biased estimate for the treatment effect and then aims to estimate the non-linear bias using RCT data.
@@ 4.3.3. Experimental Details
We adopt a multi-layer perceptron (Haykin, 1994) with 2 layers for our representation and reconstruction modules as well as the prediction heads in both stages. We tune the scale parameters in the loss functions from $51e-$ 5 to $0.1$. For the baselines with only one stage, we combine the OBS and RCT samples as the training dataset. We report both in-sample and out-of-sample results for $\sqrt{\epsilon_{\mathrm{PEHE}}}$ and $\epsilon_{\mathrm{ATE}}$ metrics in our experiments.
@@ 4.4. Performance Analysis
Table 1 shows the prediction performance with varying baselines and our methods. First, representation-based methods generally outperform generation-based methods and meta-learners, which shows the effectiveness of causal representation learning. Note that our TSPF exhibits the most competitive performance in most cases, outperforming both the representation-based and generation-based methods. More importantly, TSPF significantly outperforms another two-stage method CorNet for all metrics on the two datasets.